A cannonball is fired straight up into the air. The cannonball leaves the cannon at a speed of 63 m/s. Determine how high the cannonball goes and how long it took to reach the peak. Ignore air drag.
Answer #1

(a) v² u-agh 0= (63)-2x9.84 h - h= 202.5 m Answer (b) = u-gt 0 = 63 - 9.8+ 1 t= 6.43 s Answer